A body of mass 2kg moving with velocity of 6m/s collides with a stationary object of mass 0.5kg. If the two bodies move together after the impact, calculate their common velocity
1
Expert's answer
2019-07-01T09:29:10-0400

We can find the common velocity of the two bodies from the law of conservation of momentum:


"m_1v_{1i} + m_2v_{2i} = (m_1 + m_2) v_{com},"

here, "m_1 = 2 kg" is the mass of the body, "m_2 = 0.5 kg" is the mass of the stationary object, "v_{1i} = 6 m\/s" is the initial velocity of the body, "v_{2i} = 0" is the initial velocity of the object, "v_{com}" is the common velocity of the system of two bodies.

Then, from this equation we can calculate the common velocity of the system of two bodies:


"v_{com} = \dfrac{m_1v_{1i}}{(m_1 + m_2)} = \dfrac{2 kg \cdot 6 \dfrac{m}{s}}{(2 kg + 0.5 kg)} = 4.8 \dfrac{m}{s}."

Answer:

"v_{com} = 4.8 \dfrac{m}{s}."